A bank contains dimes and nickles there are 2 less than twice as many dimes than nickles, there are 52 coins in the bank how much money is in the bank?
Answer by stanbon(75887)

A bank contains dimes and nickles there are 2 less than twice as many dimes than nickles, there are 52 coins in the bank how much money is in the bank?
------------
Equations:
n = 2d-2
n + d = 52
---
Substitute and solve for "d":
2d-2 + d = 52
3d = 54
d = 18 (# of nickels)
------
Since n+d = 52, n = 52-18 = 34
===================================
Cheers,
Stan H.